# Data relativistic uncertainty framework for low-illumination anime scenery image enhancement

### v1

@@ 1 Introduction
In recent years, academic and industrial communities have raised a large number of attentions to the low-light enhancement (LLE) of natural images and videos [li2021low, rasheed2023comprehensive]. In contrast, the LLE task on anime scenery images has not been explored to date. Due to the artistic factor, the LLE task focuses on the perceptual and aesthetic qualities for anime scenery images, unlike mainly considering the perceptual quality for natural images. Anime images are widely shared on the Internet for many media creation purposes by general users. However, the quality of numerous anime scenery images is often degraded by different low-illumination conditions, making them unsuitable for direct use in various media creations, such as TikTok/YouTube shorts, MADmovies, posters, and animated clips. Although animation companies can re-render the original paintings to enhance the image quality, this requires laborious manual adjustments and the original paintings are not accessible to general users. Besides, the low-illumination issue can impair the robustness of other vision algorithms on anime scenery images, e.g., object detection, character segmentation and diffusion-based image editing. Hence, removing the darkness for anime scenery images in low-illumination situations is a practical and valuable task for many real-world applications.
Owing to the domain gap problem [golyadkin2025closing, zhu2021mind], current natural image low-light enhancement methods exhibit undesirable color bias and artifacts like Fig.1, when enhancing anime scenery images. As anime scenery images present distinct image characteristics from natural ones, they are out-of-distribution data [farquhar2022out] for the natural image enhancement models. Notably, available anime dataset is not existing yet for the low-illumination enhancement task. To this end, we collect anime scenery images from several open data sources and develop a method to create an essential unpaired anime scenery dataset using pseudo-data generation and a series of data processing techniques. The flexible data collection way extends the diversity of illumination conditions in unpaired dark-bright images. A certain amount of illumination data is around the mid-brightness range with large variations, making it uncertain whether the image is dark or bright. In order to quantify the illumination uncertainty and enable the model to learn with the illumination uncertainty hidden in dark/bright images, we propose a data relativistic uncertainty (DRU) framework, which is inspired by the relativistic probability concept of RGANs [jolicoeur2018relativistic]. In specific, the framework defines the data uncertainty of dark/bright image illumination as the relativistic probability (RP) between an uncertain image and the ideal dark or bright image. We consistently treat each image of the unpaired anime scenery dataset as uncertain image including those images with confident illumination to predict their RPs, which are then adopted to dynamically adjust the training loss for recalibrating optimization with the uncertainty information from dark and bright images. Since our work focuses on the data-centric learning for low-illumination anime scenery image enhancement, EnlightenGAN [jiang2021enlightengan] is chosen as the base model architecture for the significant advantages: I) multi-domain adaptability proven in real-world applications [mohamed2025integrating, wang2025unsupervised]; II) simple training scalability suitable for expanding data scale; III) lightweight unpaired training method friendly for running on most modern GPUs including 1080Ti and above. Additionally, similar to the design of most low-light enhancement methods, EnlightenGAN is unable to exploit the data uncertainty inherent in the dark/bright images.
@@ 2.1 Low-light Datasets
Natural low/normal-light image pairs are gathered by setting up one or several cameras in fixed positions and then adjusting the exposure time and ISO. This laborious collection setup constrains the diversity of data scenes and illumination conditions, making it hard to scale up the training pairs for supervised low-light image enhancement (LLIE). The commonly used paired datasets involve the first paired real-scene dataset LOL [Chen2018Retinex], the multi-exposure datasets, e.g., SCIE [cai2018learning] and SID [Chen_2018_CVPR], the photographic tone adjustment dataset broadly used in LLIE, i.e., MIT-Adobe FiveK [bychkovsky2011learning].
Unpaired datasets mainly consist of LIME [guo2016lime], NPE [wang2013naturalness], MEF [ma2015perceptual], DICM [lee2013contrast] and VV [noauthor_vasileios_nodate]. Some unpaired datasets are collected for particular application scenarios, such as ExDARK [loh2019getting] for object detection, BBD-100K [yu2020bdd100k] for autonomous driving, DARK FACE [yang2020advancing] for face recognition. Due to the tedious collection setup, these datasets suffer from a few of drawbacks including limited data scale, duplicated scene and similar illumination variation, which can lead to certain learned bias or overfitting. Several studies [wang2023teaching, xu2022transformer, matsuura2023anicropify] show that anime images have different image characteristics from natural images. However, neither paired nor unpaired datasets exist to facilitate the low-illumination anime image enhancement task. Therefore, we build the first unpaired anime scenery dataset through curating multiple data sources from other studies [jiang2023scenimefy, chen2020animegan, zhu2017unpaired, skorokhodov2021aligning]. Note that compared to natural datasets, our anime dataset exhibits more diverse scene and illumination conditions across images since it is collected in a more flexible manner than the natural datasets.
@@ 2.2 Unsupervised and Zero-Reference Enhancement
Pairwise supervision easily causes the supervised learning [lore2017llnet, cai2023retinexformer, jiang2023low] to overfit the limited training sets as reported in  [wang2024zero]. To bypass the difficult-to-scalable pairwise training, unsupervised methods enable to train with merely unpaired low/normal-light images. EnlightenGAN [jiang2021enlightengan] is the first unsupervised learning method for LLIE, which employs global-local adversarial learning, self feature preserving loss and self-regularized attention mechanism to jointly guide its model training. NeRCo [yang2023implicit] proposes a dual-closed-loop cooperative constraint strategy to train the enhancement module concurrently with controllable fitting and multi-modal supervision components in a self-supervised manner. CLIP-LIT [liang2023iterative] introduces an iterative prompt learning strategy to alternatively optimize the prompt learning framework for finer prompts and train an unsupervised enhancement network using the improved text-image similarity.
Zero-reference methods represent a higher-freedom form of unsupervised learning that does not require any paired or even unpaired data. Zero-DCE [guo2020zero] devises a series of non-reference loss functions to learn a lightweight enhancement network to estimate the parameters of a image-specific curve function, which is later refined by Zero-DCE++ [li2021learning] for better inference speed and curve estimation. Motivated by the Retinex theory, a neural architecture search strategy called RUAS [liu2021retinex] explores LLIE network from a tiny search space, while requiring an iterative process for network searching. SCI [ma2022toward] establishes a self calibrated module to converge between results of each stage, enduing stronger representation into an individual basic block. A prior-to-image framework is proposed in [wang2024zero] to train a SD enhancement network using merely normal-light data, but the physical quadruple prior is only applicable to natural images for the Kubelka–Munk theory and the model training is computationally intensive. Unsupervised and zero-reference methods often outperform the supervised counterparts in terms of the superior flexibility and generalization towards real-world applications [hu2021two, kim2022low, wang2025unsupervised]. Nevertheless, Zero-reference tasks may have difficulty in capturing reliable illumination concepts as they are learned without using any unpaired reference images. In addition, current research works are mainly model-centric. None of them attempts to explore the uncertainty information from the data with diverse illumination conditions for further data-centric learning.
@@ 2.3 Illumination Data Uncertainty
Although illumination data itself carries a certain amount of uncertainty when it is classified as low or normal light, the potential of data uncertainty remains underexplored for LLIE task. Existing related works focus on bringing uncertainty into the model weights for adaptively tackling complex low-light data. Patch-wise-based diffusion model [li2025patch] applies the uncertainty weights from each patch as a guidance for the reverse diffusion process, therefore, balancing the abrupt regions of enhanced images. Retinex-based weight uncertainty network [jin2024weight] employs Bayesian neural network to incorporate uncertainty into the enhancement model by expressing each weight as a probability distribution instead of a scalar, thereby avoiding the overconfident enhancement from fixed weights. It can be seen that their research emphasis falls on the model uncertainty for LLIE, our work rather focuses on the potential of data uncertainty for LLIE.
@@ 3.1 Unpaired Anime Data Construction Method
To rescue the dilemma of data shortage, we devise a data construction method as illustrated in Fig. 2, consisting of three stages: data aggregation, coarse data separation and refine classification of uncertain set. In the first stage, for real anime scenery data, we gather 5952 and 6196 images respectively from Scenimefy [jiang2023scenimefy] and AnimeGAN [chen2020animegan], denoted as $D_{anime}^{scenimefy}$ and $D_{anime}^{animeGAN}$. $D_{anime}^{scenimefy}$ is extracted from 9 Shinkai-style movies while $D_{anime}^{animeGAN}$ is extracted from 4 movies including Miyazaki, Shinkai and Hosoda styles. Besides, in order to expand the data scale and diversity, we use a natural scenery set $D_{natural}^{cycleGAN}$ with 6656 images from CycleGAN [zhu2017unpaired] to generate their corresponding pseudo anime scenery set $D_{pseudoanime}^{cycleGAN}$ by employing the pre-trained natural-to-anime translation model $G_{Shinkai}^{scenimefy}$ of Scenimefy. Finally, the real anime data of $D_{anime}^{scenimefy}$ and $D_{anime}^{animeGAN}$ are aggregated along with the pseudo anime data of $D_{pseudoanime}^{cycleGAN}$ as a single set $D_{anime}^{real+pseudo}$ with totally 18804 images.
@@
where $I$ is a candidate image loaded from $D_{anime}^{real+pseudo}$, while $D_{anime}^{dark}$, $D_{anime}^{bright}$ and $D_{anime}^{uncertain}$ are the resulting dark set, bright set and uncertain set separately. $mean(q(x,y))$ refers to the calculation of average brightness on the quartile patch $q(x,y)$ of candidate image $I$. $B_{low}$ and $B_{high}$ serve as the low and high boundary value for brightness. $\left\{I_{q1},...,I_{q4}\right\}$ denotes the quartile patch set obtained by evenly splitting $I$. Concretely, we use quartile patch to measure the average brightness because the quartile can be the minimal unit that contains sufficient information to influence whether an image is dark or bright, similar to the quartile application in other studies [shih2016new, okizaki2023development]. We set $B_{low}$ and $B_{high}$ to 50 and 150, based on slight adjustments to anime images through observing the statistics of natural average brightness [shi2024maco]. The average brightness calculation is written as below, where $w$ and $h$ denote the dimensions of the quartile patch.
@@ 3.2 Data Relativistic Uncertainty Framework
Motivation. In practical application scenarios, illumination is always accompanied by uncertainty. Our training datasets including trainDark and trainBright are formed by most data from uncertain set $D_{anime}^{uncertain}$, which makes the data samples in trainDark and trainBright present some data uncertainty from the ideally dark and bright image. Conventionally, the training data for EnlightenGAN [jiang2021enlightengan] merely considers dark/bright images with high-confidence brightness [Chen2018Retinex, cai2018learning], as is done in many LLIE methods [li2021low]. This practice limits the possibility of using data with comparatively uncertain brightness, which can be collected in a much easier way than with a cumbersome setup. If these methods attempt to learn directly with the dark/bright images with illumination uncertainty, they will overlook the illumination uncertainty and treat each sample with equal importance to update the loss terms. However, ignoring the illumination uncertainty from different samples will lead to unfair optimization from the updated loss terms and therefore, may cause non-smooth or under-exposed artifacts in moderately dark images (including dark images with partially bright areas) or extremely dark images after enhancement. The blue paths in Fig. 3 exhibit how the illumination uncertainty information follows the dark ($I_{d}$) and bright ($I_{b}$) images through different networks and transformation modules, ultimately leading to the undesirable update of loss terms of EnlightenGAN. By tracing each blue path back to its origin, it is found that $L_{G}^{global}$, $L_{D}^{global}$ and $L_{D}^{local}$ are all affected by the illumination uncertainty of $I_{d}$ and $I_{b}$, while $L_{G}^{local}$ is only affected by the illumination uncertainty of $I_{d}$. To sum up, how to interpretably define the illumination uncertainty, quantify the illumination uncertainty of dark/bright images, and reform the loss terms for model are the keys to learning LLIE using illumination uncertainty data.
@@
Framework Design. RaGAN has introduced a relativistic probability between the given real data and randomly sampled fake data for optimizing discriminator [jolicoeur2018relativistic]. Inspired by the relativistic probability concept of RaGAN, we propose a data relativistic uncertainty (DRU) framework to address the definition and quantification challenges of illumination uncertainty for dark/bright images, and reformulate the loss terms for EnlightenGAN to learn illumination uncertainty data. DRU framework defines the illumination uncertainty of a dark or bright image as the relativistic probability (RP) between the uncertain image and the ideally dark or bright image, which intuitively can be interpreted as the distance from the uncertain image to either the ideally dark or bright image. Although mathematically the distance between the uncertain image and the ideally dark or bright image can be calculated by existing metrics such as Euclidean distance and Cosine similarity, the ideally dark or bright image does not exist in reality and the results of these metrics cannot conform to the probability representation. Therefore, we quantify the RP between the uncertain image and the ideally dark/bright image using a learned probability network:
@@
An alternative to implement $F_{q}$ can apply centriod-based methods [ren2024deep] to compute the two centroids of dark and bright clusters by using CNN feature maps extracted from confident dark and bright images, then quantify the probabilistic distance between the uncertain image and the dark/bright centroid (roughly replacing the ideal dark/bright image). However, this is not a learnable way based on prior knowledge, so it is unable to learn about the concept approaching ideal dark/bright image.
@@
where $L_{D}^{global}$ and $L_{D}^{local}$ are respectively the global loss function for discriminator $D_{RGAN}$ [jolicoeur2018relativistic] and the local loss function for discriminator $D_{LSGAN}$ [mao2017least]. $L_{G}^{global}$ and $L_{G}^{local}$ are the global and local loss functions for generator $G$. As can be seen in Fig. 3, $I_{b}$ denotes an uncertain bright image from trainBright while $I_{d}$ is an uncertain dark image from trainDark. Accordingly, $G(I_{d})$ is the enhanced image of $I_{d}$ after passing through $G$. $\overline{I_{b}}$ and $\overline{G(I_{d})}$ refer to a randomly cropped patch from $I_{b}$ and $G(I_{d})$. Here we quantify $RP_{d}$ and $RP_{b}$ for each image of trainDark and trainBright, since even for high-confidence images, there is always tiny illumination uncertainty from the ideal dark and bright image.
Different from the primitive loss terms of EnlightenGAN [jiang2021enlightengan], the nature of our illumination-aware loss terms tends to allocate stronger gradient optimization towards dark/bright samples with lower illumination uncertainty, and conversely, weaker gradient optimization towards dark/bright samples with higher illumination uncertainty. This allows to exploit the illumination uncertainty as prior information to adaptively optimize model training based on the various illumination conditions of dark-bright image pairs. Additionally, Self Feature Preserving (SFP) losses [jiang2021enlightengan] are not affected by the illumination uncertainty of dark and bright images. Hence, we remain them to constrain the perceptual similarity between $I_{d}$ and $G(I_{d})$. Finally, EnlightenGAN aims to minimize the entire training objective:
@@
Interpretability of DRU. DRU can be understood by analogy to the wave-particle duality of light [milonni1984wave]. DRU transforms each sample into a result similar to the wave function through a probability quantizer, with two possible types: bright and dark, similar to wave states. The corresponding $RP$ represents the probability of each type. When the loss function treats it as a specific type, it’s analogous to a definite particle state. Since the probability is affected by the illumination level, the loss for a specific type can be weighted using the corresponding probability to calibrate the actual contribution of sample illumination to the loss update. Without DRU, the loss function mistakenly treats the probability of a specific type as 100%, which deviates from the actual illumination level because each sample cannot be an ideal bright or dark type. This is how illumination bias flows into the loss update during training, resulting in under- or over-enhancement of the model.

### v2

@@ 1 Introduction
In recent years, academic and industrial communities have raised a large number of attentions to the low-light enhancement (LLE) of natural images and videos [24, 39]. In contrast, the LLE task on anime scenery images has not been explored to date. Due to the artistic factor, the LLE task focuses on the perceptual and aesthetic qualities for anime scenery images, unlike mainly considering the perceptual quality for natural images. Anime scenery images are widely shared on the Internet and extensively used by general users for visual content creation. However, many such images suffer from significant quality degradation under low-illumination conditions, resulting in reduced visibility and loss of fine details, thereby limiting their direct usability.
Although animation studios can enhance image quality by re-rendering the original artworks, this process requires substantial manual effort, and the original high-quality materials are generally inaccessible to the public. Moreover, low-illumination not only degrades perceptual quality but also negatively affects the robustness of downstream vision tasks on anime scenery images, including object detection, character segmentation, and diffusion-based image editing.
Therefore, developing effective low-illumination enhancement methods tailored to anime scenery images remains a practically important yet underexplored problem with broad real-world applicability.
Current natural image low-light enhancement methods exhibit undesirable color bias and artifacts (Fig. 1) when applied to anime scenery images, due to the domain gap problem [11, 61]. Anime scenery images present distinct visual characteristics from natural images, making them effectively out-of-distribution data [10] for existing enhancement models.
Notably, no dedicated anime dataset exists for the low-illumination enhancement task. To this end, we collect anime scenery images from multiple open sources and develop a method to construct an unpaired anime scenery dataset using pseudo-data generation and a series of data processing techniques. This flexible data collection procedure extends the diversity of illumination conditions in the unpaired dark-bright images.
A certain amount of illumination data is around the mid-brightness range with large variations, making it uncertain whether the image is dark or bright. In order to quantify the illumination uncertainty and enable the model to learn with the illumination uncertainty hidden in dark/bright images, we propose a data relativistic uncertainty (DRU) framework, which is inspired by the relativistic probability concept of RGANs [20].
In specific, the framework defines the data uncertainty of dark/bright image illumination as the relativistic probability (RP) between an uncertain image and the ideal dark or bright image. We consistently treat each image of the unpaired anime scenery dataset as uncertain image including those images with confident illumination to predict their RPs, which are then adopted to dynamically adjust the training loss for recalibrating optimization with the uncertainty information from dark and bright images.
Since our work focuses on the data-centric learning for low-illumination anime scenery image enhancement, EnlightenGAN [17] is chosen as the base model architecture for the significant advantages: I) multi-domain adaptability proven in real-world applications [37, 52]; II) simple training scalability suitable for expanding data scale; III) lightweight unpaired training method friendly for running on most modern GPUs including 1080Ti and above. Additionally, similar to the design of most low-light enhancement methods, EnlightenGAN is unable to exploit the data uncertainty inherent in the dark/bright images.
@@ 2.1 Low-light Datasets
Natural low/normal-light image pairs are gathered by setting up one or several cameras in fixed positions and then adjusting the exposure time and ISO. This laborious collection setup constrains the diversity of data scenes and illumination conditions, making it hard to scale up the training pairs for supervised low-light image enhancement (LLIE). The commonly used paired datasets involve the first paired real-scene dataset LOL [54], the multi-exposure datasets, e.g., SCIE [3] and SID [5], the photographic tone adjustment dataset broadly used in LLIE, i.e., MIT-Adobe FiveK [2].
Unpaired datasets mainly consist of LIME [13], NPE [50], MEF [31], DICM [23] and VV [48]. Some unpaired datasets are collected for particular application scenarios, such as ExDARK [29] for object detection, BBD-100K [58] for autonomous driving, DARK FACE [57] for face recognition. Due to the tedious collection setup, these datasets suffer from a few of drawbacks including limited data scale, duplicated scene and similar illumination variation, which can lead to certain learned bias or overfitting. Several studies [53, 55, 34] show that anime images have different image characteristics from natural images. However, neither paired nor unpaired datasets exist to facilitate the low-illumination anime image enhancement task. Therefore, we build the first unpaired anime scenery dataset through curating multiple data sources from other studies [18, 6, 60, 46]. Note that compared to natural datasets, our anime dataset exhibits more diverse scene and illumination conditions across images since it is collected in a more flexible manner than the natural datasets.
@@ 2.2 Unsupervised and Zero-Reference Enhancement
Pairwise supervision easily causes the supervised learning [30, 4, 16] to overfit the limited training sets as reported in  [51]. To bypass the difficult-to-scalable pairwise training, unsupervised methods enable to train with merely unpaired low/normal-light images. EnlightenGAN [17] is the first unsupervised learning method for LLIE, which employs global-local adversarial learning, self feature preserving loss and self-regularized attention mechanism to jointly guide its model training. NeRCo [56] proposes a dual-closed-loop cooperative constraint strategy to train the enhancement module concurrently with controllable fitting and multi-modal supervision components in a self-supervised manner. CLIP-LIT [27] introduces an iterative prompt learning strategy to alternatively optimize the prompt learning framework for finer prompts and train an unsupervised enhancement network using the improved text-image similarity.
Zero-reference methods represent a higher-freedom form of unsupervised learning that does not require any paired or even unpaired data. Zero-DCE [12] devises a series of non-reference loss functions to learn a lightweight enhancement network to estimate the parameters of a image-specific curve function, which is later refined by Zero-DCE++ [25] for better inference speed and curve estimation. Motivated by the Retinex theory, a neural architecture search strategy called RUAS [28] explores LLIE network from a tiny search space, while requiring an iterative process for network searching.
SCI [32] establishes a self calibrated module to converge between results of each stage, enduing stronger representation into an individual basic block. A prior-to-image framework is proposed in [51] to train a SD enhancement network using merely normal-light data, but the physical quadruple prior is only applicable to natural images for the Kubelka–Munk theory and the model training is computationally intensive.
Unsupervised and zero-reference methods often outperform the supervised counterparts in terms of the superior flexibility and generalization towards real-world applications [15, 21, 52]. Nevertheless, Zero-reference tasks may have difficulty in capturing reliable illumination concepts as they are learned without using any unpaired reference images. In addition, current research works are mainly model-centric. None of them attempts to explore the uncertainty information from the data with diverse illumination conditions for further data-centric learning.
@@ 2.3 Illumination Data Uncertainty
Although illumination data itself carries a certain amount of uncertainty when it is classified as low or normal light, the potential of data uncertainty remains underexplored for LLIE task. Existing related works focus on bringing uncertainty into the model weights for adaptively tackling complex low-light data. Patch-wise-based diffusion model [26] applies the uncertainty weights from each patch as a guidance for the reverse diffusion process, therefore, balancing the abrupt regions of enhanced images. Retinex-based weight uncertainty network [19] employs Bayesian neural network to incorporate uncertainty into the enhancement model by expressing each weight as a probability distribution instead of a scalar, thereby avoiding the overconfident enhancement from fixed weights. It can be seen that their research emphasis falls on the model uncertainty for LLIE, our work rather focuses on the potential of data uncertainty for LLIE.
@@ 3.1 Unpaired Anime Data Construction Method
To rescue the dilemma of data shortage, we devise a data construction method as illustrated in Fig. 2, consisting of three stages: data aggregation, coarse data separation and refine classification of uncertain set.
In the first stage, for real anime scenery data, we gather 5952 and 6196 images respectively from Scenimefy [18] and AnimeGAN [6], denoted as $D_{anime}^{scenimefy}$ and $D_{anime}^{animeGAN}$. $D_{anime}^{scenimefy}$ is extracted from 9 Shinkai-style movies while $D_{anime}^{animeGAN}$ is extracted from 4 movies including Miyazaki, Shinkai and Hosoda styles.
To further expand the data scale and diversity, we use a natural scenery set $D_{natural}^{cycleGAN}$ with 6656 images from CycleGAN [60] to generate their corresponding pseudo anime scenery set $D_{pseudoanime}^{cycleGAN}$ by employing the pre-trained natural-to-anime translation model $G_{Shinkai}^{scenimefy}$ of Scenimefy.
Finally, the real anime data of $D_{anime}^{scenimefy}$ and $D_{anime}^{animeGAN}$ are aggregated along with the pseudo anime data of $D_{pseudoanime}^{cycleGAN}$ as a single set $D_{anime}^{real+pseudo}$ with totally 18804 images.
@@
where $I$ is a candidate image loaded from $D_{anime}^{real+pseudo}$, while $D_{anime}^{dark}$, $D_{anime}^{bright}$ and $D_{anime}^{uncertain}$ are the resulting dark set, bright set and uncertain set separately. $mean(q(x,y))$ refers to the calculation of average brightness on the quartile patch $q(x,y)$ of candidate image $I$. $B_{low}$ and $B_{high}$ serve as the low and high boundary value for brightness. $\left\{I_{q1},...,I_{q4}\right\}$ denotes the quartile patch set obtained by evenly splitting $I$.
Concretely, we use quartile patch to measure the average brightness because the quartile can be the minimal unit that contains sufficient information to influence whether an image is dark or bright, similar to the quartile application in other studies [44, 38].
We set $B_{low}$ and $B_{high}$ to 50 and 150, based on slight adjustments to anime images through observing the statistics of natural average brightness [43]. The average brightness calculation is written as below, where $w$ and $h$ denote the dimensions of the quartile patch.
@@
We adopt two sets of confident images, i.e., $D_{anime}^{dark}$ and $D_{anime}^{bright}$ to train the task which chooses ResNet18 [14] for $F_{dark/bright}$ in our experiment.
The trained classifier is used to further categorize the images of $D_{anime}^{uncertain}$ into $D_{anime}^{dark}$ or $D_{anime}^{bright}$, and a small amount of manual efforts is conducted to correct some of the medium-brightness samples with low confidence.
After this process, $D_{anime}^{dark}$ and $D_{anime}^{bright}$ contain 10303 dark images and 8501 bright images, respectively. Following an 80/20 split, $D_{anime}^{dark}$ is divided into 8240 images for trainDark and 2063 images for testDark. $D_{anime}^{bright}$ with 8501 images is directly used as trainBright.
@@ 3.2 Data Relativistic Uncertainty Framework
Motivation. In practical application scenarios, illumination is always accompanied by uncertainty. Our training datasets including trainDark and trainBright are formed by most data from uncertain set $D_{anime}^{uncertain}$, which makes the data samples in trainDark and trainBright present some data uncertainty from the ideally dark and bright image.
Conventionally, the training data for EnlightenGAN [17] merely considers dark/bright images with high-confidence brightness [54, 3], as is done in many LLIE methods [24]. This practice limits the possibility of using data with comparatively uncertain brightness, which can be collected in a much easier way than with a cumbersome setup.
If these methods attempt to learn directly with the dark/bright images with illumination uncertainty, they will overlook the illumination uncertainty and treat each sample with equal importance to update the loss terms. However, ignoring the illumination uncertainty from different samples will lead to unfair optimization from the updated loss terms and therefore, may cause non-smooth or under-exposed artifacts in moderately dark images (including dark images with partially bright areas) or extremely dark images after enhancement.
The blue paths in Fig. 3 exhibit how the illumination uncertainty information follows the dark ($I_{d}$) and bright ($I_{b}$) images through different networks and transformation modules, ultimately leading to the undesirable update of loss terms of EnlightenGAN. By tracing each blue path back to its origin, it is found that $L_{G}^{global}$, $L_{D}^{global}$ and $L_{D}^{local}$ are all affected by the illumination uncertainty of $I_{d}$ and $I_{b}$, while $L_{G}^{local}$ is only affected by the illumination uncertainty of $I_{d}$.
To sum up, how to interpretably define the illumination uncertainty, quantify the illumination uncertainty of dark/bright images, and reform the loss terms for model are the keys to learning LLIE using illumination uncertainty data.
@@
Framework Design. RaGAN has introduced a relativistic probability between the given real data and randomly sampled fake data for optimizing discriminator [20]. Inspired by the relativistic probability concept of RaGAN, we propose a data relativistic uncertainty (DRU) framework to address the definition and quantification challenges of illumination uncertainty for dark/bright images, and reformulate the loss terms for EnlightenGAN to learn illumination uncertainty data.
DRU framework defines the illumination uncertainty of a dark or bright image as the relativistic probability (RP) between the uncertain image and the ideally dark or bright image, which intuitively can be interpreted as the distance from the uncertain image to either the ideally dark or bright image.
Although mathematically the distance between the uncertain image and the ideally dark or bright image can be calculated by existing metrics such as Euclidean distance and Cosine similarity, the ideally dark or bright image does not exist in reality and the results of these metrics cannot conform to the probability representation. Therefore, we quantify the RP between the uncertain image and the ideally dark/bright image using a learned probability network:
@@
An alternative to implement $F_{q}$ can apply centriod-based methods [40] to compute the two centroids of dark and bright clusters by using CNN feature maps extracted from confident dark and bright images, then quantify the probabilistic distance between the uncertain image and the dark/bright centroid (roughly replacing the ideal dark/bright image). However, this is not a learnable way based on prior knowledge, so it is unable to learn about the concept approaching ideal dark/bright image.
@@
where $L_{D}^{global}$ and $L_{D}^{local}$ are respectively the global loss function for discriminator $D_{RGAN}$ [20] and the local loss function for discriminator $D_{LSGAN}$ [33]. $L_{G}^{global}$ and $L_{G}^{local}$ are the global and local loss functions for generator $G$. As can be seen in Fig. 3, $I_{b}$ denotes an uncertain bright image from trainBright while $I_{d}$ is an uncertain dark image from trainDark. Accordingly, $G(I_{d})$ is the enhanced image of $I_{d}$ after passing through $G$. $\overline{I_{b}}$ and $\overline{G(I_{d})}$ refer to a randomly cropped patch from $I_{b}$ and $G(I_{d})$. Here we quantify $RP_{d}$ and $RP_{b}$ for each image of trainDark and trainBright, since even for high-confidence images, there is always tiny illumination uncertainty from the ideal dark and bright image.
Different from the primitive loss terms of EnlightenGAN [17], the nature of our illumination-aware loss terms tends to allocate stronger gradient optimization towards dark/bright samples with lower illumination uncertainty, and conversely, weaker gradient optimization towards dark/bright samples with higher illumination uncertainty. This allows to exploit the illumination uncertainty as prior information to adaptively optimize model training based on the various illumination conditions of dark-bright image pairs. Additionally, Self Feature Preserving (SFP) losses [17] are not affected by the illumination uncertainty of dark and bright images. Hence, we remain them to constrain the perceptual similarity between $I_{d}$ and $G(I_{d})$. Finally, EnlightenGAN aims to minimize the entire training objective:
@@
Interpretability of DRU. DRU can be understood by analogy to the wave-particle duality of light [35]. Each sample is transformed into a representation similar to a wave function via a probability quantizer, with two possible states: bright and dark. The corresponding $RP$ indicates the probability of each state.
When the loss function treats a sample as a specific state, this is analogous to observing a definite particle state. Since the probability reflects the illumination level, the loss for each state can be weighted by its corresponding probability, thus calibrating each sample’s contribution to the loss update. Without DRU, the loss function assumes a 100% probability for a specific state, ignoring the actual illumination uncertainty. This leads to illumination bias propagating through the loss updates, causing under- or over-enhancement in the resulting model.